A jeweler is using 10 lbs. of an alloy that is 2% robium. He has another alloy which is 5% robium. He wishes to combine these mixtures to form a mixture that is 3% robium. How many pounds of the second alloy should be added to the first to get a mixture that is 3% robium.
I really need help, so I would ask for step by step instructions please.
Thank you.

Answer by stanbon(75887)

A jeweler is using 10 lbs. of an alloy that is 2% robium. He has another alloy which is 5% robium. He wishes to combine these mixtures to form a mixture that is 3% robium. How many pounds of the second alloy should be added to the first to get a mixture that is 3% robium.
-------------------
2% alloy DATA:
Amount = 10 lbs ; amount of robium= 0.02(10)=0.2 lbs
--------------
5% alloy DATA:
Amount= x lbs ; amount of robium=0.05x lbs.
------------
Mixture DATA:
Amount = x+10 lbs ; amount of robium = 0.3(x+10)=(0.03x+0.3) lbs
----------------
EQUATION:
robium + robium = robium
0.2 +0.05x = 0.03x+0.3
0.02x=0.1
x=5 lbs (amount of 5% alloy that must be added)
Cheers,
Stan H.